OpenStudy (anonymous):

Can someone please solve this and show work? Directions - Factor completely. -28y^2 + 7t^2

OpenStudy (anonymous):

Rewrite it as: \[7t^2-28y^2=7(t^2-4y^2)=7(t+2y)(t-2y)\] Because: \[(a^2-b^2)=(a-b)(a+b)\] Called a difference of squares.